OpenStudy (anonymous):

The circumference of a pizza is approximately 44 inches. Estimate the area of the pizza.. A. 154 in2 B. 14 in2 C. 7 in2 D. 53 in2

jhonyy9 (jhonyy9):

the circumference formula is 2pir so from this you can calcule r ,so after you will get r so will be very easy to calcule the area

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

so than 2pir=44 pir=22 r=22/pi area=pir^2 area=pi(22/pi)^2 area=((22)^2)/pi

jhonyy9 (jhonyy9):

so 154 in^2 is answer right,correct

jhonyy9 (jhonyy9):

A. is right

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

Yes thanks(:

jhonyy9 (jhonyy9):

yw good luck bye

jhonyy9 (jhonyy9):

so if you can understanding this calcules you will can solving your next exercise easy